Lovell Computer Parts Inc. is in the process of setting a selling price on a new component it has just designed and developed. The following cost estimates for this new component have been provided by the accounting department for a budgeted volume of 50,000 units.

Per Unit Total
Direct materials $48
Direct labor $21
Variable manufacturing overhead $17
Fixed manufacturing overhead $750,000
Variable selling and administrative expenses $22
Fixed selling and administrative expenses $350,000


Lovell Computer Parts management requests that the total cost per unit be used in cost-plus pricing its products. On this particular product, management also directs that the target price be set to provide a 20% return on investment (ROI) on invested assets of $1,000,000.

Compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 20% on this new component. (Round markup percentage to 2 decimal places, e.g. 10.50%.)

Markup percentage %
Target selling price $

eTextbook and Media

Assuming that the volume is 40,000 units, compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 20% on this new component. (Round answers to 2 decimal places, e.g. 10.50% or 10.50.)

Markup percentage %
Target selling price $
Answer #1
1
Direct materials 48
Direct labor 21
Variable manufacturing overhead 17
Fixed manufacturing overhead 15 =750000/50000
Variable selling and administrative expenses 22
Fixed selling and administrative expenses 7 =350000/50000
Total cost per unit 130
Desired ROI per unit 4 =(1000000*20%)/50000
Desired ROI per unit 4
Divide by Total cost per unit 130
Markup Percentage 3.08%
Desired ROI per unit 4
Add: Total cost per unit 130
Target Selling price 134
2
Direct materials 48.00
Direct labor 21.00
Variable manufacturing overhead 17.00
Fixed manufacturing overhead 18.75 =750000/40000
Variable selling and administrative expenses 22.00
Fixed selling and administrative expenses 8.75 =350000/40000
Total cost per unit 135.50
Desired ROI per unit 5.00 =(1000000*20%)/40000
Desired ROI per unit 5.00
Divide by Total cost per unit 135.50
Markup Percentage 3.69%
Desired ROI per unit 5.00
Add: Total cost per unit 135.50
Target Selling price 140.50